OpenStudy (abbie):

solve for a. 25a+ab=5b

OpenStudy (anonymous):

wat is A

OpenStudy (abbie):

? O.o

OpenStudy (***[isuru]***):

25a + ab = 5b a(25 + b) = 5b a = (5b)/ (25 + b) i think this is what u r looking for :)

OpenStudy (anonymous):

\[25a+ab=5b\] \[a(25+b)=5b\] \[a=5b/(25+b)\]

OpenStudy (abbie):

thank you mustafa :)

OpenStudy (anonymous):

umm the solution to this equation could not be determined, because for two variables 2 equations are needed but if you want to rearrange then @mustafa2014 & @***[ISURU]*** both did right !

OpenStudy (anonymous):

yw :)